OpenStudy (anonymous):

find the exact value of the arc length intercepted by a central angle of 105 degrees in a circle with a radius of 9 centimeters

jhonyy9 (jhonyy9):

do you know what is formula of circumference of a cyrcle ?

jhonyy9 (jhonyy9):

C = 2*pi*r

jhonyy9 (jhonyy9):

yes ?

OpenStudy (jdoe0001):

\(\bf \textit{Arc's Length}=\cfrac{radius\cdot angle\cdot \pi}{180}\)